Reports 1-1 of 1 Clear search Modify search
masayuki.nakano - 2:08 Wednesday 25 September 2019 (10687) Print this report
ETMY MN Coil driver inspection
[Okutomi, Nakano]

We found harmonics in the output signal of the coil driver. The details are as follows:


  • We rebooted the k1ey1 PC, but it did not make a drastic difference. Here, the k1ey1 could not be turned on for a while. The detail is reported by Yamamoto-san.
  • Then, we went to the Yend to check the coil drivers. I connected the V monitor to the empty ADC port (ADC1_CH24~27). I put the sinusoidal signal into K1:VIS-ETMY_MN_LOCK_L_EXC as the previous measurement (freq. of 0.2 Hz, amp. of 10000 cnts). 
  • The spectrum of the VMON signal is shown in the attached file. You can see many harmonics in the signal, and the highest frequency of harmonics is about 50 Hz.
  • The harmonics have a different amplitude in each channel. This can cause the imbalanced force to each coil and excite yaw motion at the harmonic frequencies.
  • I forget to disable the drive align between L and P. Therefore, there is the signal on the V3 channel. But again, here are the harmonic signals and this can cause pitch motion at the harmonic frequencies.
  • I also check the other coil driver, which is for the MN and the IM vertical coils. I put the excitation signal in K1:VIS-ETMY_MN_LOCK_P_EXC with same amplitude and frequency as the privious excitation, which means only the MN V2 is excited. The second figure shows that V2 channel also has harmonic signals.
  • I forgot to check the output signal of the AI chassis. So still, it can be the problem of the DAC or the AI filter. Tomorrow, we should check that.
  • We also did the same excitation measurement at ITMY, and it did not seem to have such a HUM problem. It will be reported by Okutomi. It would be worthful to check the coil driver output signal of the ITMY.
  • I did not check the consistency between these harmonic signals and the actual motion of the mirror. We also need to check it to conclude this caused the strange behavior of the suspension.



  • To connect the V MON port to the ADC, we need an adapter made by Shimode-san. I left it in the end station with the Dsub-cable in the rack for the AA/AI chassis. You should connect the adapter side to the VMON port.
  • One of the new HP coil drivers was brought to the end station. I also left it. We can try to replace it, but if you do that, please remember that the output filter of the coils need to be adjusted to have same actuator efficiency with the current coil driver.
Images attached to this report
Comments to this report:
koki.okutomi - 2:48 Wednesday 25 September 2019 (10689) Print this report
  • I tried to reproduce the same nonlinear behavior in the ITMY payload.
    • Even with a large amplitude of excitation such as > 10000 cnts in MN_COILOUTF_H, I could not observe any strange response of the TM OpLevs.
    • I attach some figures for examples of the TM OpLev signals with MNL excitation in ITMY.
Images attached to this comment
Search Help